OpenStudy (anonymous):

Graph the functions and identify the vertical asymptote. Please help and explain so I can do it myself?

OpenStudy (anonymous):

1. y = 1 / x - 4 minus 2

OpenStudy (anonymous):

2. y = 1 / x + 2 plus 4

OpenStudy (anonymous):

just get the equations to the form x/a + y/b =1 and try to get the value of b. It will be the vertical asymptote I hope.

OpenStudy (anonymous):

Oh, okay! Could you walk me through it as I do it?

OpenStudy (anonymous):

I think im too over-qualified to help you on this....

OpenStudy (anonymous):

Please? Anything helps. I just really need a good explainer. (:

OpenStudy (austinl):

Do you have a graphing calculator?

OpenStudy (anonymous):

No :(

OpenStudy (austinl):

hmmmm.... Ok, for this the graph would look like this. The vertical asmyptote is located at x=4 The horizontal asmyptote is located at y=-2